# Multiple Instance Learning for Glioma Diagnosis using Hematoxylin and Eosin stained Whole Slide Images: An Indian Cohort Study

### 24 Feb 2024 v1

@@ 7 Conclusion
Our study showcases the potential of AI in enhancing brain tumor classification and IHC biomarker detection through advanced machine learning models. Employing a ResNet-50 architecture pre-trained with the Barlow Twin SSL approach and utilizing DTFD for feature aggregation, we achieved significant improvements in diagnostic precision across various histopathological datasets. This approach validates traditional diagnostic criteria and unveils novel morphological insights, enriching clinical understanding and patient care. Incorporating AI into pathology heralds a new era of diagnostics, offering profound implications for personalized treatment and patient outcomes, demonstrating AI’s critical role in revolutionizing medical diagnostics and therapeutic strategies.

### 08 Mar 2024 v2

@@ 7 Conclusion
Our study demonstrated how MIL-based algorithms can significantly improve brain tumor diagnosis by classifying subtypes, grading, and IHC biomarkers. We achieved significant improvements in diagnostic precision across various histopathological datasets employing a ResNet-50 architecture pre-trained with the Barlow Twin SSL approach and utilizing DTFD for feature aggregation. Experimentation with diverse configurations of feature extraction and aggregation led to this setup that outperformed current state-of-the-art solutions. Our methodology focuses on using H&E stained slides exclusively, aiming to improve diagnostics in settings where IHC staining may be logistically or financially challenging. We plan to expand our research to include a wider range of feature extractors and aggregators, exploring or developing optimal combinations that could further improve diagnostic performance.
8 Compliance with Ethical Standards
Procedures in studies with human participants adhered to ethical standards set by institutional (NIMS-IEC) and/or national research committees (ICMR).